A particle moves along x-axis according to expression x=3t²-2t where x is in metre and t in seconds.



a) What is the displacement of the particles in the time interval



t=0 to t=1



t=1 to t=4




b) Calculate the average velocity for the time intervals in (a)



c) Find the instantenous velocity for a particle at 3 seconds

1
Expert's answer
2021-12-07T20:11:37-0500

Given:

"x(t)=3t\u00b2-2t"


a) the displacement of the particles


"d_{01}=x(1)-x(0)=3*1^2-2*1-0=1\:\rm m"

"d_{14}=x(4)-x(1)\\=(3*4^2-2*4)-(3*1^2-2*1)=39\:\rm m"

b) the average velocity of the particle

"v_{01}=\frac{d_{01}}{\Delta t}=\rm\frac{1\: m}{1\: s}=1\: m/s"

"v_{14}=\frac{d_{14}}{\Delta t}=\rm\frac{39\: m}{3\: s}=13\: m/s"

c) the instantaneous velocity


"v(t)=x'(t)=(3t^2-2t)'=6t-2"

at the instant "t=3\:\rm s"

"v(3)=6*3-2=16\:\rm m/s"